OpenStudy (anonymous):

A courtyard is shaped like a square with 250-ft-long sides. What is the distance from one corner of the courtyard to the opposite corner?

OpenStudy (anonymous):

pythagorean theorem, where your answer is going to be the length of the hypotenuse (c). 250^2 +250^2 = c^2

OpenStudy (anonymous):

353.6 ft

OpenStudy (anonymous):

d = sqrt ( 2 * 250^2) = 250 sqrt(2) = 353.6